# The role of charm and unflavored mesons in prompt atmospheric lepton fluxes

### v1

@@ 4.2 Flux ratios
The ratios of the angle-averaged fluxes $\phi_{\bar{\nu}_{\mu}}/\phi_{\nu_{\mu}}]$ are shown in the right panel of fig. 10. The value of this ratio at $E_{\nu}=10^{3}$ GeV, amounting to $\sim 0.6$, is less than unity because of the same charge excess of $\pi^{+}$ relative to $\pi^{-}$ that yields $\phi_{\mu^{+}}/\phi_{\mu^{-}}]>1$ at the same energy (see discussion in the previous paragraph). The ratio $\phi_{\bar{\nu}_{\mu}}/\phi_{\nu_{\mu}}]$ increases to $\sim 1$ at the energy where the prompt neutrino flux dominates. The light unflavored mesons do not contribute to the neutrino flux so the neutrino flux does not depend on $\alpha_{\rm unfl}$. The ratio curves differ by $\sim 10\%$ for $E_{\nu}=10^{6}$ GeV. This will have little impact on the ($\nu_{\mu}+\bar{\nu}_{\mu}$) interaction rates, especially since the neutrino and antineutrino cross sections differ by less than 5% at this energy Weigel:2024gzh .
@@
Figure 11 shows the ratio of $\phi_{\mu^{+}+\mu^{-}}/\phi_{\nu_{\mu}+\bar{\nu}_{\mu}}$ as a function of energy for different choices for $w_{c}^{\rm intr}$ and $\alpha_{\rm unfl}$. The two lower curves use as input $\alpha_{\rm unfl}=1$ and either $w_{c}^{\rm intr}=0$ or $w_{c}^{\rm intr}=\penalty 10000\ 4.65\times 10^{-4}$. The black curve for $w_{c}^{\rm intr}=0$ tends to a ratio of 2 in the figure since the unflavored and perturbative charm contributions to the muon flux are approximately equal, and the perturbative charm contributions to the neutrino and muon fluxes are nearly equal. For $E\gtrsim 10^{6}$ GeV, including intrinsic charm decreases the ratio by $\sim 18\%$, since intrinsic charm is relatively more important for the neutrino flux than for the muon flux. At the energies where ratios with $\alpha_{\rm unfl}=1$ shown in fig. 11 differ, the astrophysical neutrino flux already dominates the atmospheric neutrino flux. Extracting the ratio in this energy regime will be difficult. The upper two curves in fig. 11 include the best-fit values of $\alpha_{\rm unfl}$ for $w_{c}^{\rm intr}=0$ and $4.65\times 10^{-4}$. The largest values of the ratio $\phi_{\mu^{+}+\mu^{-}}/\phi_{\nu_{\mu}+\bar{\nu}_{\mu}}$ are obtained when the unflavored meson contribution to the prompt $\mu^{+}+\mu^{-}$ flux is increased with no increase in the prompt $\nu_{\mu}+\bar{\nu}_{\mu}$ flux, namely when $w_{c}^{\rm intr}=0$ and $\alpha_{\rm unfl}=4.80$.
@@
Figure 11: Ratio of muon flux to muon neutrino flux with and without intrinsic charm angle averaged for $\theta_{\rm zen}\leq 60^{\circ}$. The red (black) curve has $\alpha_{\rm unfl}=1$, and $w_{c}^{\rm intr}=4.65\times 10^{-4}$ ($w_{c}^{\rm intr}=0$). The green (blue) curve has $w_{c}^{\rm intr}=4.65\times 10^{-4},\ \alpha_{\rm unfl}=4.33$ ($w_{c}^{\rm intr}=0,\ \alpha_{\rm unfl}=4.8$).
@@
Figure 12: Ratio of muon flux per solid angle $\phi_{\mu^{+}+\mu^{-}}(\theta)$ at zenith angles of $0^{\circ}$ and $60^{\circ}$ to the angle-averaged $\theta_{\rm zen}\leq 60^{\circ}$ muon flux per solid angle $\phi_{\mu^{+}+\mu^{-}}$ without any intrinsic charm (labeled MCEq flux), with $w_{c}^{intr}=4.65\times 10^{-4}$ and $\alpha_{\rm unfl}=1$ (with intrinsic charm), and with $w_{c}^{intr}=4.65\times 10^{-4}$ and $\alpha_{\rm unfl}=4.33$ (with scaling of unflavored meson contributions).
@@
Finally, in fig. 12 we show the ratio of the muon flux per solid angle at zenith angles $\theta_{\rm zen}=0^{\circ}$ and $60^{\circ}$ relative to the angle-averaged muon flux. For the flux dominated by pion decays, the flux scales as $\sim 1/\cos\theta$ for the zenith angle range considered here Klimushin:2000cy . With this functional form, the flux ratios are $\phi_{\mu^{+}+\mu^{-}}(0^{\circ})/\phi_{\mu^{+}+\mu^{-}}(\leq 60^{\circ})=0.72$ and $\phi_{\mu^{+}+\mu^{-}}(60^{\circ})/\phi_{\mu^{+}+\mu^{-}}(\leq 60^{\circ})=1.44$, consistent with the curves in fig. 12 for $E_{\mu}\sim 10^{4}$ GeV. The dominance of prompt contributions, whether from perturbative charm, intrinsic charm or light unflavored mesons, leads to an isotropic distribution and a unit ratio at high energy, independent of angle. Figure 12 shows the trend towards prompt dominance as the ratio curves move closer to one. The transition to prompt-dominated atmospheric muon flux as a function of energy is more rapid with $\alpha_{\rm unfl}=4.33$ than with just intrinsic charm or with the default MCEq.
@@ 4.3 Uncertainties
There are a number of uncertainties in the MCEq evaluation of the atmospheric leptons fluxes. We discuss them assuming $w_{c}^{\rm intr}=0$. In our MCEq computations, we do not include any seasonal variations in the atmospheric muon and muon neutrino fluxes. The IceCube collaboration has measured a 3.94.63.9-4.6 % seasonal modulation in the lepton flux for energies of 125 GeV - 10 TeV IceCube:2025wtk , a small variation compared to other uncertainties and mostly relevant to the conventional atmospheric lepton contributions.
@@
A potentially larger uncertainty comes from the hadronic interaction modeling. Within the MCEq framework, a range of Monte Carlo event generators for hadronic interactions are available: besides Sibyll-2.3c, EPOS-LHC Pierog:2013ria , QGSJETII04 Ostapchenko:2010vb and DPMJETIII306 Roesler:2000he can also be used for predictions. The three panels of fig.13 show ratios of fluxes using these three models to our default result using Sibyll-2.3c for the angle-average ($\theta_{\rm zen}\leq 60^{\circ}$) for $\mu^{+}+\mu^{-}$. The upper left panel shows the ratios of the prompt fluxes from light unflavored mesons. DPMJETIII306 has a larger proportion of prompt muons from light unflavored mesons than Sibyll-2.3c does for $E_{\mu}\gtrsim 4\times 10^{5}$ GeV, and almost a factor of two larger for $E_{\mu}=10^{6}$ GeV. The EPOS-LHC angle-averaged muon flux from light unflavored mesons nearly equals that from Sibyll-2.3c. The QGSJETII04 results for muons from light unflavored mesons are $\sim 20\%$ of those from Sibyll-2.3c.
@@
With Sibyll-2.3c, we showed that the prompt atmospheric flux of muons from light unflavored mesons is nearly equal to the prompt atmospheric muon flux from charm. The upper right panel of fig. 13 shows the $\mu^{+}+\mu^{-}$ flux ratios for all prompt atmospheric muons (unflavored plus charm). Neither QGSJETII04 nor EPOS-LHC include charm production, so the ratios are consequently reduced. The prompt atmospheric muon flux from DPMJETIII306 does include charm contributions. Including both unflavored mesons and charm, the ratio is $\sim 1.8$ for $E_{\mu}=10^{6}$ GeV and $\sim 0.85$ for $E_{\mu}=10^{4}$ GeV. The lower panel of fig. 13 shows the ratios of the total $\mu^{+}+\mu^{-}$ flux. In the lower energy range of the figure, the ratios are within $\sim 10\%$, while at high energies where the prompt flux dominates (or has no charm contribution as in the cases of QGSJETII04 and EPOS-LHC), the curves deviate. Comparing the muon fluxes of DPMJETIII306 and Sibyll-2.3c, the total flux ratio is $\sim 0.9$ for $E_{\mu}=10^{5}$ GeV, and $\sim 1.3$ for $E_{\mu}=10^{6}$ GeV. A best-fit rescaling of the DPMJETIII306 unflavored meson contribution gives $\alpha_{\rm unfl}=4.84$ (without the inclusion of an intrinsic charm component). Even though the DPMJETIII306 muon flux for $E_{\mu}=10^{6}$ GeV is larger, the fit is dominated by the muon data with smaller errors, $E_{\mu}\lesssim 3\times 10^{5}$ GeV, where the DPMJETIII306 flux is slightly smaller than the Sibyll-2.3c flux, so the value of $\alpha_{\rm unfl}$ is slightly larger for the DPMJETIII306 muon flux than for the Sibyll-2.3c flux.
@@
In fig. 15, we show the ratio of the fluxes evaluated with the H4a Gaisser:2013bla (sometimes called H3p, for instance in ref. IceCube:2016umi ) cosmic ray model relative to the H3a model for muons (left) and muon neutrinos (right). Both H3a and H4a are implemented in MCEq as 3-Peters cycle Peters:1961mxb compositions with five mass groups. However, H4a assumes a composition that becomes proton dominated above the ankle, whereas H3a retains a mixed composition. The H4a model predicts a larger high-energy lepton flux, leading to an enhancement of approximately 35% for muons and 40% for muon neutrinos at $E=5\times 10^{6}$ GeV. However, this results in only modest changes to the fitted intrinsic charm and unflavored scaling parameters. For the Regge1 intrinsic charm model, the best fit values shift to $w_{c}^{\rm intr}=4.35\times 10^{-4},\ \alpha_{\rm unfl}=1$ and $w_{c}^{\rm intr}=0,\ \alpha_{\rm unfl}=3.85$ when using H4a instead of H3a, compared to the values reported in table 2.
@@
In ref. Yanez:2023lsy , Yañez and Fedynitch tune a data-driven model (daemonflux) that includes hadronic interaction and cosmic ray flux uncertainties. The tune is calibrated to atmospheric muon flux data, primarily below $E_{\mu}=10^{4}$ GeV. They compare the resulting neutrino flux from daemonflux with the atmospheric neutrino fluxes from Sibyll-2.3d, which yields nearly identical inclusive atmospheric lepton fluxes as Sibyll-2.3c Riehn:2019jet . The Sibyll-2.3d muon neutrino flux is within approximately $\pm 5\%$ of the daemonflux muon neutrino flux for $E_{\nu_{\mu}}=10^{4}-10^{6}$ GeV. There is a little more variation in the electron neutrino flux, with the Sibyll-2.3d electron neutrino flux approximately $15\%$ higher for $E_{\nu_{e}}=10^{4}$ GeV, and closer to 5% higher for $E_{\nu_{e}}=10^{5}$ GeV. Our conclusion is that using a data-driven model such as daemonflux would not change our estimates about uncertainties in the atmospheric muon flux predictions, so there remains an apparent deficit in atmospheric muon flux predictions at high energy.
@@
Different intrinsic charm models can, in principle, yield widely varying atmospheric flux predictions. We have assumed a functional form for the differential cross section for charm hadron production via the intrinsic charm mechanism in nucleon-air interactions according to eq. 11. The resulting contributions from intrinsic-charm-produced $D^{-},\ \bar{D}^{0}$, and $\Lambda_{c}$ to the muon and muon neutrino flux have essentially the same shape. The primary effect of changing the fragmentation model is a shift in the overall normalization of the intrinsic charm contribution, which is reflected in the fitted value of $w_{c}^{\rm intr}$, as shown in tables 1 and 2. The reason the fitted $w_{c}^{\rm intr}$ values differ among the models is that the Regge models produces a larger fraction of $\Lambda_{c}$ baryons at higher energies compared to $D$ mesons, whereas the HLM model yields almost equal distributions of energetic $D$ and $\Lambda_{c}$ hadrons. Since $D$ mesons have higher branching ratios to muon and muon neutrino than $\Lambda_{c}$, and the leptons from $D$ meson decays can emerge with higher energy fractions than for $\Lambda_{c}$ decays (see figs. 18 and 19 in the Appendix), the HLM model gives a relatively larger intrinsic charm contribution to the muon and muon neutrino fluxes, so the best fit normalization $w_{c}^{\rm intr}$ is smaller for HLM fragmentation compared to the Regge models. In other words, the physical differences between the models manifest primarily as a change in the fitted intrinsic charm normalization, rather than in the shape of the resulting lepton spectra. Overall, whether $w_{c}^{\rm intr}$ is fitted to the muon flux or the muon neutrino flux, all three valence-like fragmentation models for intrinsic charm give values for $w_{c}^{\rm intr}$ that are roughly comparable among each other.
5 Discussion and Conclusions
IceCube’s measurements of the $\mu^{+}+\mu^{-}$ atmospheric flux at high energies suggest a missing prompt component. We have focused on fluxes at zenith angles $\theta\leq 60^{\circ}$ to avoid issues with potential mis-reconstruction of muon events in the experimental data IceCube:2015wro . Since the prompt $\nu_{\mu}+\bar{\nu}_{\mu}$ and $\mu^{+}+\mu^{-}$ fluxes from charm are nearly equal, an increase of the prompt $\mu^{+}+\mu^{-}$ flux from charm to improve agreement with the high-energy muon data (which would correspond to taking $w_{c}^{\rm intr}=3.87\times 10^{-3}$) can violate the upper bounds on the prompt $\nu_{\mu}+\bar{\nu}_{\mu}$ flux. Indeed, in addition, such a high value of $w_{c}^{\rm intr}$ would lead to an excess of very forward production of $\Lambda_{c}^{+}$ in low-energy collisions since we have assumed that $w_{c}^{\rm intr}$ is energy independent. The value of $w_{c}^{\rm intr}=4.65\times 10^{-4}$ is sufficiently small that the intrinsic charm cross section is less than 5% of the perturbative cross section for fixed-target $pp$ production of $c\bar{c}$ at $E_{p}=10^{5}$ GeV and less than $\sim 1\%$ for $E_{p}\gtrsim 10^{6}$ GeV, well within the next-to-leading order perturbative Nason:1987xz ; Bhattacharya:2016jce ; Garzelli:2015psa and measurement PHENIX:2006tli ; HERA-B:2007rfd ; ALICE:2012inj ; STAR:2012nbd ; ATLAS:2015igt ; LHCb:2013xam uncertainties for open charm production. However, this smaller value of $w_{c}^{\rm intr}$ still underpredicts the atmospheric $\mu^{+}+\mu^{-}$ flux at high energy, as shown in fig. 7.

### v2

@@ 4.2 Flux ratios
The ratios of the angle-averaged fluxes $\phi_{\bar{\nu}_{\mu}}/\phi_{\nu_{\mu}}]$ are shown in the right panel of fig. 11. The value of this ratio at $E_{\nu}=10^{3}$ GeV, amounting to $\sim 0.6$, is less than unity because of the same charge excess of $\pi^{+}$ relative to $\pi^{-}$ that yields $\phi_{\mu^{+}}/\phi_{\mu^{-}}]>1$ at the same energy (see discussion in the previous paragraph). The ratio $\phi_{\bar{\nu}_{\mu}}/\phi_{\nu_{\mu}}]$ increases to $\sim 1$ at the energy where the prompt neutrino flux dominates. The light unflavored mesons do not contribute to the neutrino flux so the neutrino flux does not depend on $\alpha_{\rm unfl}$. The ratio curves differ by $\sim 10\%$ for $E_{\nu}=10^{6}$ GeV. This will have little impact on the ($\nu_{\mu}+\bar{\nu}_{\mu}$) interaction rates, especially since the neutrino and antineutrino cross sections differ by less than 5% at this energy Weigel et al. (2025).
@@
Figure 12 shows the ratio of $\phi_{\mu^{+}+\mu^{-}}/\phi_{\nu_{\mu}+\bar{\nu}_{\mu}}$ as a function of energy for different choices for $w_{c}^{\rm intr}$ and $\alpha_{\rm unfl}$. The two lower curves use as input $\alpha_{\rm unfl}=1$ and either $w_{c}^{\rm intr}=0$ or $w_{c}^{\rm intr}=\penalty 10000\ 1.01\times 10^{-3}$. The black curve for $w_{c}^{\rm intr}=0$ tends to a ratio of 2 in the figure since the unflavored and perturbative charm contributions to the muon flux are approximately equal, and the perturbative charm contributions to the neutrino and muon fluxes are nearly equal. For $E\gtrsim 10^{6}$ GeV, including intrinsic charm decreases the ratio by $\sim 18\%$, since intrinsic charm is relatively more important for the neutrino flux than for the muon flux. At the energies where ratios with $\alpha_{\rm unfl}=1$ shown in fig. 12 differ, the astrophysical neutrino flux already dominates the atmospheric neutrino flux. Extracting the ratio in this energy regime will be difficult. The upper two curves in fig. 12 include the best-fit values of $\alpha_{\rm unfl}$ for $w_{c}^{\rm intr}=0$ and $1.01\times 10^{-3}$. The largest values of the ratio $\phi_{\mu^{+}+\mu^{-}}/\phi_{\nu_{\mu}+\bar{\nu}_{\mu}}$ are obtained when the unflavored meson contribution to the prompt $\mu^{+}+\mu^{-}$ flux is increased with no increase in the prompt $\nu_{\mu}+\bar{\nu}_{\mu}$ flux, namely when $w_{c}^{\rm intr}=0$ and $\alpha_{\rm unfl}=4.86$.
@@
Figure 12: Ratio of muon flux to muon neutrino flux with and without intrinsic charm angle averaged for $\theta_{\rm zen}\leq 60^{\circ}$. The red (black) curve has $\alpha_{\rm unfl}=1$, and $w_{c}^{\rm intr}=1.01\times 10^{-3}$ ($w_{c}^{\rm intr}=0$). The green (blue) curve has $w_{c}^{\rm intr}=1.01\times 10^{-3},\ \alpha_{\rm unfl}=3.84$ ($w_{c}^{\rm intr}=0,\ \alpha_{\rm unfl}=4.86$).
@@
Figure 13: Ratio of muon flux per solid angle $\phi_{\mu^{+}+\mu^{-}}(\theta)$ at zenith angles of $0^{\circ}$ and $60^{\circ}$ to the angle-averaged $\theta_{\rm zen}\leq 60^{\circ}$ muon flux per solid angle $\phi_{\mu^{+}+\mu^{-}}$ without any intrinsic charm (labeled MCEq flux), with $w_{c}^{intr}=1.01\times 10^{-3}$ and $\alpha_{\rm unfl}=1$ (with intrinsic charm), and with $w_{c}^{intr}=1.01\times 10^{-3}$ and $\alpha_{\rm unfl}=3.84$ (with scaling of unflavored meson contributions).
@@
Finally, in fig. 13 we show the ratio of the muon flux per solid angle at zenith angles $\theta_{\rm zen}=0^{\circ}$ and $60^{\circ}$ relative to the angle-averaged muon flux. For the flux dominated by pion decays, the flux scales as $\sim 1/\cos\theta$ for the zenith angle range considered here Klimushin et al. (2001). With this functional form, the flux ratios are $\phi_{\mu^{+}+\mu^{-}}(0^{\circ})/\phi_{\mu^{+}+\mu^{-}}(\leq 60^{\circ})=0.72$ and $\phi_{\mu^{+}+\mu^{-}}(60^{\circ})/\phi_{\mu^{+}+\mu^{-}}(\leq 60^{\circ})=1.44$, consistent with the curves in fig. 13 for $E_{\mu}\sim 10^{4}$ GeV. The dominance of prompt contributions, whether from perturbative charm, intrinsic charm or light unflavored mesons, leads to an isotropic distribution and a unit ratio at high energy, independent of angle. Figure 13 shows the trend towards prompt dominance as the ratio curves move closer to one. The transition to prompt-dominated atmospheric muon flux as a function of energy is more rapid with $\alpha_{\rm unfl}=3.84$ than with just intrinsic charm or with the default MCEq.
@@ 4.3 Uncertainties
There are a number of uncertainties in the MCEq evaluation of the atmospheric leptons fluxes. We discuss them assuming $w_{c}^{\rm intr}=0$. In our MCEq computations, we do not include any seasonal variations in the atmospheric muon and muon neutrino fluxes. The IceCube collaboration has measured a 3.94.63.9-4.6 % seasonal modulation in the lepton flux for energies of 125 GeV - 10 TeV Abbasi and others (2025b), a small variation compared to other uncertainties and mostly relevant to the conventional atmospheric lepton contributions.
@@
A potentially larger uncertainty comes from the hadronic interaction modeling. Within the MCEq framework, a range of Monte Carlo event generators for hadronic interactions are available: besides Sibyll-2.3c, EPOS-LHC Pierog et al. (2015), QGSJETII04 Ostapchenko (2011) and DPMJETIII306 Roesler et al. (2000) can also be used for predictions. The three panels of fig.14 show ratios of fluxes using these three models to our default result using Sibyll-2.3c for the angle-average ($\theta_{\rm zen}\leq 60^{\circ}$) for $\mu^{+}+\mu^{-}$. The upper left panel shows the ratios of the prompt fluxes from light unflavored mesons. DPMJETIII306 has a larger proportion of prompt muons from light unflavored mesons than Sibyll-2.3c does for $E_{\mu}\gtrsim 4\times 10^{5}$ GeV, and almost a factor of two larger for $E_{\mu}=10^{6}$ GeV. The EPOS-LHC angle-averaged muon flux from light unflavored mesons nearly equals that from Sibyll-2.3c. The QGSJETII04 results for muons from light unflavored mesons are $\sim 20\%$ of those from Sibyll-2.3c.
@@
With Sibyll-2.3c, we showed that the prompt atmospheric flux of muons from light unflavored mesons is nearly equal to the prompt atmospheric muon flux from charm. The upper right panel of fig. 14 shows the $\mu^{+}+\mu^{-}$ flux ratios for all prompt atmospheric muons (unflavored plus charm). Neither QGSJETII04 nor EPOS-LHC include charm production, so the ratios are consequently reduced. The prompt atmospheric muon flux from DPMJETIII306 does include charm contributions. Including both unflavored mesons and charm, the ratio is $\sim 1.8$ for $E_{\mu}=10^{6}$ GeV and $\sim 0.85$ for $E_{\mu}=10^{4}$ GeV. The lower panel of fig. 14 shows the ratios of the total $\mu^{+}+\mu^{-}$ flux. In the lower energy range of the figure, the ratios are within $\sim 10\%$, while at high energies where the prompt flux dominates (or has no charm contribution as in the cases of QGSJETII04 and EPOS-LHC), the curves deviate. Comparing the muon fluxes of DPMJETIII306 and Sibyll-2.3c, the total flux ratio is $\sim 0.9$ for $E_{\mu}=10^{5}$ GeV, and $\sim 1.3$ for $E_{\mu}=10^{6}$ GeV. A best-fit rescaling of the DPMJETIII306 unflavored meson contribution gives $\alpha_{\rm unfl}=4.89$ (without the inclusion of an intrinsic charm component). Even though the DPMJETIII306 muon flux for $E_{\mu}=10^{6}$ GeV is larger, the fit is dominated by the muon data with smaller errors, $E_{\mu}\lesssim 3\times 10^{5}$ GeV, where the DPMJETIII306 flux is slightly smaller than the Sibyll-2.3c flux, so the value of $\alpha_{\rm unfl}$ is slightly larger for the DPMJETIII306 muon flux than for the Sibyll-2.3c flux.
@@
In fig. 16, we show the ratio of the fluxes evaluated with the H4a Gaisser et al. (2013) (sometimes called H3p, for instance in ref. Aartsen and others (2016b)) cosmic ray model relative to the H3a model for muons (left) and muon neutrinos (right). Both H3a and H4a are implemented in MCEq as 3-Peters cycle Peters (1961) compositions with five mass groups. However, H4a assumes a composition that becomes proton dominated above the ankle, whereas H3a retains a mixed composition. The H4a model predicts a larger high-energy lepton flux, leading to an enhancement of approximately 35% for muons and 40% for muon neutrinos at $E=5\times 10^{6}$ GeV, however, this results in only modest changes to the fitted intrinsic charm and unflavored scaling parameters. For the Regge1 intrinsic charm model, the best fit values shift to $w_{c}^{\rm intr}=0.94\times 10^{-4},\ \alpha_{\rm unfl}=1$ and $w_{c}^{\rm intr}=0,\ \alpha_{\rm unfl}=3.91$ when using H4a instead of H3a, compared to the values reported in table 2.
By exploiting their new measurements of the cosmic ray energy spectrum and mean logarithmic mass Lv et al. (2025), together with other high-energy cosmic-ray data, the LHAASO collaboration models the cosmic ray spectrum and composition with three galactic and two extragalactic components. Their model has a harder spectrum and lighter composition in the PeV cosmic ray energy range. This new cosmic ray flux model can yield energy dependent variations in the prompt atmospheric lepton fluxes from charm by $20\%\ -\ +50\%$ for the interval $E=10^{4}-10^{6}$ GeV. In any case, the increase around $10^{6}$ GeV is not sufficient to fully account for the measured muon flux. Further investigations of the impact of the spectrum and composition of galactic cosmic rays are ongoing Das et al. (2026).
@@
There are uncertainties in modeling hadronic interactions of cosmic rays with air targets. In ref. Yañez and Fedynitch (2023), Yañez and Fedynitch tune a data-driven model (daemonflux) that includes hadronic interaction and cosmic ray flux uncertainties. The tune is calibrated to atmospheric muon flux data, primarily below $E_{\mu}=10^{4}$ GeV. They compare the resulting neutrino flux from daemonflux with the atmospheric neutrino fluxes from Sibyll-2.3d, which yields nearly identical inclusive atmospheric lepton fluxes as Sibyll-2.3c Riehn et al. (2020). The Sibyll-2.3d muon neutrino flux is within approximately $\pm 5\%$ of the daemonflux muon neutrino flux for $E_{\nu_{\mu}}=10^{4}-10^{6}$ GeV. There is a little more variation in the electron neutrino flux, with the Sibyll-2.3d electron neutrino flux approximately $15\%$ higher for $E_{\nu_{e}}=10^{4}$ GeV, and closer to 5% higher for $E_{\nu_{e}}=10^{5}$ GeV. Our conclusion is that using a data-driven model such as daemonflux would not change our estimates about uncertainties in the atmospheric muon flux predictions, so there remains an apparent deficit in atmospheric muon flux predictions at high energy.
@@
Different intrinsic charm models can, in principle, yield widely varying atmospheric flux predictions. We have assumed a functional form for the differential cross section for charm hadron production via the intrinsic charm mechanism in nucleon-air interactions according to eq. 11. The resulting contributions from intrinsic-charm-produced $D^{-},\ \bar{D}^{0}$, and $\Lambda_{c}$ to the muon and muon neutrino flux have essentially the same shape. The primary effect of changing the fragmentation model is a shift in the overall normalization of the intrinsic charm contribution, which is reflected in the fitted value of $w_{c}^{\rm intr}$, as shown in tables 1 and 2. The reason the fitted $w_{c}^{\rm intr}$ values differ among the models is that the Regge models produces a larger fraction of $\Lambda_{c}$ baryons at higher energies compared to $D$ mesons, whereas the HLM model yields almost equal distributions of energetic $D$ and $\Lambda_{c}$ hadrons. Since $D$ mesons have higher branching ratios to muon and muon neutrino than $\Lambda_{c}$, and the leptons from $D$ meson decays can emerge with higher energy fractions than for $\Lambda_{c}$ decays (see figs. 19 and 20 in the Appendix), the HLM model gives a relatively larger intrinsic charm contribution to the muon and muon neutrino fluxes, so the best fit normalization $w_{c}^{\rm intr}$ is smaller for HLM fragmentation compared to the Regge models. In other words, the physical differences between the models manifest primarily as a change in the fitted intrinsic charm normalization, rather than in the shape of the resulting lepton spectra. Overall, whether $w_{c}^{\rm intr}$ is fitted to the muon flux or the muon neutrino flux, all three valence-like fragmentation models for intrinsic charm give values for $w_{c}^{\rm intr}$ that are roughly comparable among each other.
5 Discussion